IDVI Features

Features of IDVI include the following: (Click on each feature to show more details.)

1. Displays arbitrary DVI files.

(a) Does not require any particular macro package. Works with plain TEX, L^{A}TEX, L^{A}TEX 2ε, etc.

(b) Handles unusual fonts and styles.

(c) Displays equations and tables:

∑_{n=1}^{∞} \frac{1}{n^{2}} = \frac{π^{2}}{6} \begin{aligned}& 1 & a & b \\ 1 & 1 & a & b \\ a & a & b & 1 \\ b & b & 1 & a \end{aligned}

2. Displays encapsulated postscript images , even in the magnifier window.

3. Displays colored text.

4. Provides interaction through hypertext links.

(a) The HyperTEX specials are interpreted, to provide both internal and external links.

(b) Special idvi: URLs can be used to control navigation from within a document. You can provide controls to increase or decrease magnification, for example.

(c) Although links which span pages are not allowed, multiline links do work.

5. Provides more interaction through text hiding and outlines.

(a) If you can see this text, you are using the outline feature.

(b) Nested outlines work as expected.

I’m not sure what to use them for in a document such as this one, however.

(c) Outlines can be controlled by links placed anywhere on the page. For example, the sections of this document which can be hidden describe: display of arbitrary DVI files, hypertext links, text hiding, applets, and portability.

6. Allows fully interactive documents, through the use of embedded Java applets .

7. Runs on any platform which supports the Java language.

(a) The DVI Viewer software runs inside any web browser which supports Java.

(b) The idvi document preparation software runs on any Unix machine which supports Java. (The Java part of the software receives assistance from some shell scripts, and so will run not correctly on Windows or Macintosh machines without some more work.)

However, there are presently some real limitations to the use of IDVI:

1. All fonts must be downloaded from the web server which provides the dvi file. This can be slow, particularly if the server is heavily loaded.

2. The DVI Viewer code itself is somewhat large, and is slow to download.

3. IDVI does not support virtual fonts.

Please see the IDVI user’s guide for more information about using the features demonstrated above.

The L^{A}TEX source for this document may also be of interest.